Refer the table below. Our original information was that the probability of High demand was .6. We collected the following new information about what happened to demand when there was an improving or not improving economy. Answer the below question.

Calculate the probability for high demand and the improving economy?

Improving Economy Not improving economy
Given high demand 0.75
Given Low demand 0.45

Submit

Answer format: Number: Round to: 2 decimal places.
Answer #1

P(Improving Economy/High Demand)=0.75

P(Not Improving Economy/High Demand)=0.25

P(Improving Economy/Low Demand)=0.55

P(Not Improving Economy/Low Demand)=0.45

P(Prior Probability of high Demand)=0.6

P(Prior Probability of Low Demand)=1-0.6=0.4

P(High Demand and improving economy)=P(improving economy/High Demand )*P(Prior probability of High demand)

P(High Demand and improving economy)=0.75*0.6 =0.45

Answer= 0.45